JEF meeting, October 11, 2024
From GlueXWiki
Connection
Meeting time and location:
- 9:30 am EST (JLab time) remote only
- Connection information
Agenda
Preparation of parts for ECAL -- Sasha
- - PCBs with 5 (12 x 4) and 5 (6 x 8) connectors are ready for two distribution boards. One more board will be ready during next week.
- - Boxes for HV boards have been ordered. The lead time about 3 weeks.
- - Communicated with Vince P. (JLab's machine shop) and Tim. Ordered 2 temporary distribution boards, if the delivery of the real ones is delayed
Installation of the FCAL2 detector & integration to GlueX -- Sasha, All
- Installation of the light distribution board of the ECAL LMS [3]
- - Fine-tune the light source. Slightly modified the design, ordered new LEDs
- Implementing ECAL into the GlueX DAQ [4]
- - rocecal3 crate was fixed [5] (initial DAQ implementation [6])
- - implementing flash ADC scalers [7]
- - translation table for the ECAL (finalize/check translation DB, generate xml file, import to the CCDB) - done [8]
- - implementing and verifying ECAL LMS trigger [9]
- - implementing and verifying the ECAL physics trigger. First event reconstructed using the GlueX software [10]
- - Installation of scintillating paddles for cosmic tests [11]
- - Start implementing VTP trigger modules. Requires modification of firmware. The trigger crates may be unavailable for 1 - 1.5 months.
- ECAL HV interlock check [12]
- Cable map of the ECAL LG modules after installing ECAL [13]
- Reinstalling cables on the ECAL LV distribution board [14][15]
- Safety briefings - Vlad, Sasha
- Detector installation page
- FCAL2 Engineering Meetings
- Remaining installation tasks
- - Connecting cooling pipes to the ECAL fan (done by tech group)
- - Installing dark room ceiling and insulation inside the room, required for cooling test (almost complete)
- - Test cooling system (in progress, almost complete)
- - Installing HV distribution boards, start switching the detector ON (waiting for HV boxes)
Restoring PS test setup [16]
ECAL commissioning -- Sasha, All
Minutes
Attendees: Bill, Igal, Sasha, Vlad, Zhikun, Liping, Ilya, S.T.
- HV distribution boards in house, being populated with connectors, but box to house them has still not been delivered
- Box has been ordered and PR signed, expect delivery in ~3 weeks
- In the meantime planning for temporary box
- LV distribution boards lowered, need to re-cable
- LMS
- Tuning of light source underway
- Use more LEDs and/or increase power
- Modify to use narrower aperature LEDs -- Vlad Popov
- Some fibers need repair
- Reference PMT: may need additional magnetic shielding, alpha source needs to be coordinated with RadCon
- Tuning of light source underway
- Integration of ECAL into DAQ
- ECAL LMS trigger now working
- GTP trigger with small number of channels implemented
- Working on VTP-based trigger, needs firmware update
- Recorded waveform for one PMT --> translation table and low-level reconstruction library working!
- Translation table for lead glass region implemented
- Dark room almost complete
- Checked interlock system, humidity/temperature sensors not included yet
- Dehumidier(s) will be installed
- ECAL commissioning plan being fleshed out
- Expect will need ~1 month
- Start with 2V setting on flash-ADCs, but may need to go down to 1V, need data to decide
- Need to minimize gain variation to ~5% level for trigger
- ECAL has been added to ESAD